OpenStudy (maheshmeghwal9):

Solve: - \[\huge{\color{purple}{(\frac{x+i}{x-i})^n=1.}}\]

OpenStudy (anonymous):

Go by rationalizing the denominator by multiplying and dividing by x + i

OpenStudy (maheshmeghwal9):

ok wait a min. plz

OpenStudy (anonymous):

Take your time..

OpenStudy (maheshmeghwal9):

I gt \[(\frac{x^2-1+2x i}{x^2+1})^n=1\]

OpenStudy (anonymous):

You can take both the sides nth root.. can you??

OpenStudy (maheshmeghwal9):

i have just studied that so i can't:(

OpenStudy (anonymous):

\[a^n = 1\] \[\sqrt[n]{a^n} = \sqrt[n]{1}\] \[\huge \color{green} {a^{n \times \frac{1}{n}} = (1)^{\frac{1}n}}\] Till here you got it?

OpenStudy (maheshmeghwal9):

ya

OpenStudy (anonymous):

I think I am going wrong wait a minute..

OpenStudy (maheshmeghwal9):

ok

OpenStudy (anonymous):

see anything power 0 is 1.. so n should be 0 there..

OpenStudy (maheshmeghwal9):

ok

OpenStudy (maheshmeghwal9):

BUT i m nt getting the answer @waterineyes

OpenStudy (anonymous):

Yes but tell me for which variable will we solve is it for n or x??

OpenStudy (maheshmeghwal9):

We have to solve for x & Its answer is \[\cot \frac{k \pi}{n}\]where k= 1,2,3.....n

OpenStudy (anonymous):

See till here you got it: \[(\frac{x^2 - 1 + 2i.x}{x^2 + 1})^n = 1\]

OpenStudy (maheshmeghwal9):

ok then?

OpenStudy (anonymous):

\[(\frac{x^2 - 1}{x^2 + 1} + \frac{2x}{x^2 + 1}i)^n = 1\] Ok??

OpenStudy (maheshmeghwal9):

ok:)

OpenStudy (anonymous):

Now we have to make a little substitution: Put: \[\frac{x^2 - 1}{x^2 + 1} = sinx\] now can you find cosx from here?? \[cosx = \sqrt{1 - \sin^2x}\] Find cosx from here..

OpenStudy (maheshmeghwal9):

ya i can do

OpenStudy (anonymous):

Solving or I should solve this??

OpenStudy (maheshmeghwal9):

solving........

OpenStudy (anonymous):

Take your time..

OpenStudy (maheshmeghwal9):

I gt \[(\sin x + i \cos x)^n=1\]\[\implies [ \cos (\frac{\pi}{2}-x)+ i \sin (\frac{\pi}{2}-x)]^n=1.\]now wt should i do?

OpenStudy (anonymous):

Do you know about the DE MOIVER'S THEOREM??

OpenStudy (maheshmeghwal9):

of course so i should apply it here.

OpenStudy (anonymous):

Yes you should..

OpenStudy (maheshmeghwal9):

\[[\cos n(\frac{\pi}{2}-x)+i \sin n(\frac{\pi}{2}-x)]=1\]

OpenStudy (maheshmeghwal9):

now?

OpenStudy (anonymous):

See, complex number comprises of real and imaginary part.. Since on RHS there is no imaginary part so equate real part of LHS with real part of RHS..

OpenStudy (maheshmeghwal9):

bt then i m getting \[x=\frac{\pi}{2}.\]

OpenStudy (anonymous):

No no.. We have to use here general formula: Do not solve just equate and show me what it has become now.

OpenStudy (maheshmeghwal9):

ok \[\cos n(\frac{\pi}{2}-x)=1\]\[i \sin n(\frac{\pi}{2}-x)=0i=0.\]

OpenStudy (anonymous):

Wait a minute..

OpenStudy (maheshmeghwal9):

ok:)

OpenStudy (anonymous):

I am not getting the solution that you want but:: The General solution of: \[\cos \theta = \cos x\] is: \[\theta = 2k \pi \pm x\] where k belongs to set of integer.. So, here \[cosn(\frac{\pi}{2} - x) = \cos0\] \[n(\frac{\pi}{2} - x) = 2k \pi \pm 0\] From here calculate x but the answer is not coming right..

OpenStudy (maheshmeghwal9):

ok np i gt ur solution. but i think u r right:)

OpenStudy (maheshmeghwal9):

thanx for ur gr8 help:)

OpenStudy (anonymous):

Thanks and medal for what?? Answer is not coming right that you want..

OpenStudy (maheshmeghwal9):

medal for gr8 help:)

OpenStudy (anonymous):

We are very close to it..

OpenStudy (maheshmeghwal9):

ya:)